A professor finds that for 9 randomly selected students the time needed to complete an assignment...

  1. A professor finds that for 9 randomly selected students the time needed to complete an assignment had a sample mean of 60 minutes and a sample standard deviation of 18 minutes. The 98% confidence interval estimate for the population mean is:

    a.

    50.99 - 77.01

    b.

    42.62 - 77.38

    c.

    59.00 - 69.00

    d.

    41.37 - 65.63

QUESTION 26

  1. After estimating the interval in the previous question, we can interpret it as:

    a.

    We are 98% confident that the true mean falls into the estimated interval.

    b.

    We are 98% confident that the sample mean falls into the estimated interval.

    c.

    We are 98% confident that the estimated interval contains the true mean.

    d.

    We are 98% confident that the estimated interval contains the sample mean.
Answer #1

Solution :

t /2,df = 2.896

Margin of error = E = t/2,df * (s /n)

= 2.896 * (18 / 9)

Margin of error = E = 17.38

The 98% confidence interval estimate of the population mean is,

- E < < + E

60 - 17.38 < < 60 + 17.38

42.62 < < 77.38

42.62 - 77.38

b)

We are 98% confident that the estimated interval contains the true mean.